# The quantum criticality of the Standard Model and the hierarchy problem

### v1

@@ 6 Phase diagram deformations and new physics
In the bottom panels, we vary $c_{3}$ (left panel) and $c_{4}$ (right). Positive deformations of these parameters lower the position of the horizontal partial fixed point. In fact, this triggers the stabilization of the quartic coupling towards the UV and ultimately lowers the associated Landau pole. These deformations are interesting because they improve or even cure the metastability of the Higgs sector. Also, they reduce the fine-tuning in the direction of the Higgs parameters. This minimization occurs at two instances. First, at the maximum values of the deformations shown, $c_{3}\simeq 0.2$ and $c_{4}\simeq 0.04$, which correspond to the scenarios in which the trajectories displaying the SM IR values have a stable shape and $\bar{\mu}^{2}_{\Lambda_{\textrm{UV}}}=0$ (again entering the same class as that discussed in Goertz:2023pvn ). The second instance occurs for $c_{3}\simeq-0.1$ and $c_{4}\simeq-0.005$, and corresponds to the case where the quartic coupling vanishes at the UV scale. However, it is important to note that the sensitivity with respect to other (marginal) coupling directions remains rather constant. As the dominant sensitivity is still sourced in the Yukawa direction, which is rather insensitive to these deformations, the total BG measure remains unchanged and no fine-tuning is relaxed. However, we emphasize that in more realistic model embeddings, where specific combinations of $c_{i}$ are required, the dependence on deformations could differ, potentially revealing directions that reduce the overall fine-tuning. This is in particular also true for scenarios that couple the new physics differently to the different SM sectors, which calls for further investigation.
@@
In summary, we observe that it is not easy to find more natural configurations in theory space via the corrections considered here and in the presence of a common UV scale, other than by reducing the canonical dimension of the Higgs mass operator Aoki:2012xs . The dominant sensitivity in the marginal coupling directions arises as a direct consequence of the mixing between the relevant direction and the marginal couplings. The most promising pathway appears to be the simplest one: introducing new physics that effectively lowers the dimension of the Higgs mass operator. In fact, this has been explored in the context of self-organisation Bornholdt:1992up ; Gies:2025pqv ; Wetterich:2016uxm . Also, this direction includes models where the Higgs field operators can be formulated in a more fundamental way from higher-order fermionic interactions; see e.g. Kaplan:1983fs ; Kaplan:1983sm ; Dugan:1984hq ; Hill:2002ap ; Agashe:2004rs ; Cacciapaglia:2020kgq ; Cacciapaglia2022a for Composite Higgs models and Ahmed:2023qsm for similar recent proposals which complete Little Higgs models in a UV-safe way, as well as Goertz:2023nii ; Goertz:2024dnz for treatments within a first-principles approach. Other interesting implementations in this direction could be achieved by coupling an elementary Higgs to a strong sector or by considering a large multiplicity of new dark fields Gies:2025pqv . All these directions benefit from a non-perturbative treatment of the new-physics scenario, which will be explored in more detail in an independent work.
@@ 7 Conclusions
Naturalness is often used as a guiding principle in the formulation of SM extensions, yet its meaning, closely tied to subtle quantum corrections in the Higgs mass, remains interpretation dependent. In this work we have revisited the notion of fine-tuning present in the hierarchy problem, from the viewpoint of critical phenomena. For this, we have studied the phase structure of the SM class of theories using the functional Renormalization Group (fRG), a Wilsonian and mass-dependent RG scheme that offers a distinct perspective to the common perturbative approaches. Within this framework, we strengthen the interrelation between the hierarchy problem and the near-quantum criticality of the SM.
@@
We began by reviewing the RG flow of the Higgs mass and its implications on the theory’s phase diagram and special trajectories. The IR regularization intrinsic to the fRG, together with its ability to account for power-law running, allowed us to track the evolution of the effective Higgs potential along the RG-flow, akin to a finite-temperature system. This feature enabled a detailed analysis of the phase diagram, including the identification of the critical surface and associated distinct trajectories (e.g. scale invariant), in a intuitive and graphical way.
@@
We investigated the properties of the quantum phase transition that separates theories with and without a Higgs mechanism. The associated critical behavior is characterized by universal quantities such as critical exponents, which we linked directly to standard measures of fine-tuning. Notably, we emphasized the conceptual connection between the proximity of the SM to criticality and its sensitivity to UV physics —two phenomena that, while distinct, share a common physical origin in the presence of new physics. This connection proves particularly useful in the context of naturalness, as it allows one to formulate the degree of tuning in a physical and scheme-independent manner. We have stressed that, within this finite RG approach—where no divergences are encountered—the IR–UV sensitivity can be expressed in terms of well-defined and universally derivable quantities. This stands in contrast with some interpretations in other finite schemes Mooij:2024rys ; Mooij:2021lbc ; Mooij:2021ojy .
@@
The low-energy effective theory viewpoint adopted here provides a general framework for studying the sensitivity of IR observables to unknown UV completions in an agnostic way. Within this setting, we identified two distinct sources of fine-tuning. First, the well-known tuning due to the separation between the electroweak scale and that of new physics (mimicked by a UV cutoff), and second, an additional source stemming from the theory’s proximity to the critical surface. While the former leads to the standard quadratic sensitivity, the latter generates even stronger dependencies, particularly for theories near the quantum phase transition. In fact, some mild heavy new-physics deformations can eliminate the Higgs mechanism entirely, rendering radically different IR spectra. This latter tuning dimension is, in principle, accessible in any scheme and could open new avenues for model building in the context of the little hierarchy problem.
@@
To illustrate the practical utility of our framework, we presented an example where we search for low-energy extensions only coupled to the Higgs sector which reduce the sensitivity of IR observables to heavy new physics. Within this simplistic analysis, we recovered well-known mechanisms involving large anomalous dimensions for the Higgs field, thus demonstrating that such scenarios can be naturally embedded and analyzed within our approach.

### v2

@@ 6 Phase diagram deformations and new physics
In the bottom panels, we vary $c_{3}$ (left panel) and $c_{4}$ (right). Positive deformations of these parameters lower the position of the horizontal partial fixed point. In fact, this triggers the stabilization of the quartic coupling towards the UV and ultimately lowers the associated Landau pole. These deformations are interesting because they improve or even cure the metastability of the Higgs sector. Also, they reduce the fine-tuning in the direction of the Higgs parameters. This minimization occurs at two instances. First, at the maximum values of the deformations shown, $c_{3}\simeq 0.2$ and $c_{4}\simeq 0.04$, which correspond to the scenarios in which the trajectories displaying the SM IR values have a stable shape and $\bar{\mu}^{2}_{\Lambda_{\textrm{UV}}}=0$ (again entering the same class as that discussed in Goertz:2023pvn ). The second instance occurs for $c_{3}\simeq-0.1$ and $c_{4}\simeq-0.005$, and corresponds to the case where the quartic coupling vanishes at the UV scale. However, it is important to note that the sensitivity with respect to other (marginal) coupling directions remains rather constant. As the dominant sensitivity is still sourced in the Yukawa direction, which is rather insensitive to these deformations, the total BG measure remains unchanged and no fine-tuning is relaxed. However, we emphasize that in more realistic model embeddings, where specific combinations of $c_{i}$ are required, the dependence on deformations could differ, potentially revealing directions that reduce the overall fine-tuning. This is in particular also true for scenarios that couple the new physics differently to the different SM sectors, which calls for further investigation. Especially in the context of the little hierarchy problem, where a mild improvement is sufficient, more natural models could be found.
@@
In summary, we observe that it is not easy to find more natural configurations in theory space via the simple corrections considered here and in the presence of a common UV scale, other than by reducing the canonical dimension of the Higgs mass operator Aoki:2012xs . The dominant sensitivity in the marginal coupling directions arises as a direct consequence of the mixing between the relevant direction and the marginal couplings. The most promising pathway appears to be the simplest one: introducing new physics that effectively lowers the dimension of the Higgs mass operator. In fact, this has been explored in the context of self-organisation Bornholdt:1992up ; Gies:2025pqv ; Wetterich:2016uxm . Also, this direction includes models where the Higgs field operators can be formulated in a more fundamental way from higher-order fermionic interactions; see e.g. Kaplan:1983fs ; Kaplan:1983sm ; Dugan:1984hq ; Hill:2002ap ; Agashe:2004rs ; Cacciapaglia:2020kgq ; Cacciapaglia2022a for Composite Higgs models and Ahmed:2023qsm for similar recent proposals which complete Little Higgs models in a UV-safe way, as well as Goertz:2023nii ; Goertz:2024dnz for treatments within a first-principles approach. Other interesting implementations in this direction could be achieved by coupling an elementary Higgs to a strong sector or by considering a large multiplicity of new dark fields Gies:2025pqv . All these directions benefit from a non-perturbative treatment of the new-physics scenario, which will be explored in more detail in an independent work.
@@ 7 Conclusions
Naturalness of the electroweak scale provides one of the main guiding principles in the search for a more fundamental realization of nature beyond the Standard Model (SM), and it is inherently rooted in the quantum corrections to the Higgs mass. In order to mitigate the large quantum corrections which come along with a severe fine-tuning, new physics is generally expected to appear at scales of order TeV. The absence of new physics at the LHC suggests that the scale of new physics should be higher, leading to the so-called little hierarchy problem. As a result, proposed solutions to the big hierarchy problem face additional constraints, since the generic new physics scale needs to be raised while still remaining natural (or very different mechanisms need to be considered). Most approaches along these lines rely on mechanisms that postpone the full impact of the hierarchy problem. There might or might not be a fundamental link between such mechanisms and the physics invoked to address the full hierarchy problem, though such a connection might be a hint towards a consistent solution.
To add to the understanding of these pressing questions, in this work we revisited the notion of fine-tuning present in the hierarchy problem from the viewpoint of critical phenomena. For this, we have studied the phase structure of the SM class of theories using the functional Renormalization Group (fRG), a Wilsonian and mass-dependent renormalization group scheme that offers a distinct perspective to the common perturbative approaches. The developed framework allowed deepening the understanding of the hierarchy problem by emphasizing that SM-like theories with a new scale are not only quadratically fine-tuned in the conventional sense, but also near quantum critical. We stressed that, extrapolating the low energy description up to some large new physics scale not only imposes conditions on quadratic corrections, but logarithmic contributions must also be considered in the tuning, which is relevant for UV completions of the SM. In particular, we showed that the near criticality implies an enhanced tuning in the neighborhood of the SM theory space that should be acknowledged in naturalness studies, especially regarding the little hierarchy problem.
@@
In this paper we first reviewed the RG flow of the Higgs mass and its implications on the theory’s phase diagram and special trajectories. The IR regularization intrinsic to the fRG, together with its ability to account for power-law running, allowed us to track the evolution of the effective Higgs potential along the RG-flow, akin to a finite-temperature system. This feature enabled a detailed analysis of the phase diagram, including the identification of the critical surface and associated distinct trajectories (e.g. the scale invariant one), in a intuitive and graphical way.
@@
We then investigated the properties of the quantum phase transition that separates theories with and without a Higgs mechanism. The associated critical behavior is characterized by universal quantities such as critical exponents, which we linked directly to standard measures of fine-tuning. Notably, we emphasized the conceptual connection between the proximity of the SM to criticality and its sensitivity to UV physics — two phenomena that, while distinct, share a common physical origin in the presence of new physics. This connection proves particularly useful in the context of naturalness, as it allows one to formulate the degree of tuning in a physical and scheme-independent manner. We have stressed that, within this finite RG approach — where no divergences are encountered — the IR–UV sensitivity can be expressed in terms of well-defined and universally derivable quantities. This stands in contrast with some interpretations in other finite schemes Mooij:2024rys ; Mooij:2021lbc ; Mooij:2021ojy .
@@
The low-energy effective theory viewpoint adopted here provides a general framework for studying the sensitivity of IR observables to unknown UV completions in an agnostic way. Within this setting, we identified two distinct sources of fine-tuning. First, the well-known tuning due to the separation between the electroweak scale and that of new physics (mimicked by a UV cutoff), and second, an additional source stemming from the theory’s proximity to the critical surface. While the former leads to the standard quadratic sensitivity, the latter generates even stronger dependencies, particularly for theories near the quantum phase transition. In fact, some mild heavy new-physics deformations can eliminate the Higgs mechanism entirely, rendering radically different IR spectra. This latter tuning dimension is, in principle, accessible in any scheme and could open new avenues for model building, also in the context of solving the little hierarchy problem.
@@
To illustrate the practical utility of our framework, we presented an example where we search for low-energy extensions only coupled to the Higgs sector which reduce the sensitivity of IR observables to heavy new physics. Within this simplistic analysis, we also recovered well-known mechanisms involving large anomalous dimensions for the Higgs field, thus demonstrating that such scenarios can be naturally embedded and analyzed within our approach.